OpenStudy (anonymous):

Help what is 25(d-12)=4 &explain how to solve it

OpenStudy (anonymous):

25d-300=4 collect the like terms to one side: 25d=300+4 can u carry on to the answer now?

OpenStudy (anonymous):

No I'm still confused

OpenStudy (anonymous):

what specifically confuses you?

OpenStudy (anonymous):

Expand the brackets. (25x d) - (25 x-12)=4 This would give you: 25d-300= 4 Move -300 to the other side by doing the opposite of -300 which is +300. 25d-300+300= 4+300 (-300 +300=0 (so it cancels out and is left with 25d=300+4)) Simplify: 25d=304 divide both sides by25: 25d/25 = 304/25 d=12.16